OpenStudy (anonymous):

If an object is thrown directly upwards (initial speed 3m/s) and the only other force acting on it is gravity (9.8m/s) what is the maximum height it will reach?

OpenStudy (mrnood):

I assume you know the equations for constant acceleration v=u+at v^2=u^2+2as s=ut+0.5at^2 In this case you know that u = 3m/s, its maximum height will be achieved when v=0 you know a = -9.8m/s^2 (it is NEGATIVE because the velocity is being measured as positive in UP direction, but acceleration of gravity is positive in a downward direction) so you know u,v,a Choose the equation above that will let you solve s

OpenStudy (anonymous):

\[h = Vit + \frac{ 1 }{ 2 }g t ^{2}\]

OpenStudy (anonymous):

for upward movement of body g = -g

OpenStudy (mrnood):

@IhteshamMalik That is not the best equation - t is not known

OpenStudy (anonymous):

\[2gh = Vf ^{2} - Vi ^{2}\]

OpenStudy (mrnood):

Now - it would be good if the OP solved it for himself.....

OpenStudy (anonymous):

Vi = 3 m/s g = -g Vf = 0 so, hmm next OP can solve himself...